A ball moving at 1 m.s-1 strikes another ball twice the mass which is initially at rest.

The lighter ball bounces back at 0.5 m.s-1, its path making an angle of 30° with its

original path. Calculate the velocity of the heavier ball after impact.


1
Expert's answer
2021-04-12T06:54:34-0400

"m\vec{v}=m\vec{v'}+2m\vec{u},"

"mv=-mv'cos30\u00b0+2mu,"

"u=\frac{v+v'cos30°}{2}=\frac{1+0.5\cdot 0.87}{2}=0.72~\frac ms."